Margarite wants to take a trip to japan in 8 years. she knows she will need about $4200.how much should she deposit into an account now that pays 3,5% interest comppound daily,in order to meet her goal?round to the nearest dollar.
Answer by Phillips54(20)

x/4200=3.5/8
4200*3.5=14700
14700/8=1837.5
x=1837.5
she started out with $1837.50